OpenStudy (anonymous):

if v=8 find the value of 3.v

OpenStudy (karatechopper):

is tht 3 time v? or 3 point v?

OpenStudy (saifoo.khan):

24

OpenStudy (anonymous):

times

OpenStudy (karatechopper):

24

OpenStudy (anonymous):

3*v = 3*8 = 24

OpenStudy (anonymous):

24